# Quantum Bootstrap Approach to a Non-Relativistic Potential for Quarkonium systems

### v1

@@ IV Bootstrap Implementation and Results for Charmonium and Bottomonium
The results show the effectiveness of the quantum bootstrap method. For the ground states ($l=0$), the predicted masses for $J/\psi$ and $\Upsilon$ are accurate, with deviations of less than 2 MeV and 6 MeV, corresponding to an accuracy of over 99.98(94)%. The predictions for the P-wave states ($l=1$) are also strong, with deviations of less than 7 MeV. This slight deviation is expected, as this model calculates the spin-averaged mass and does not include spin-dependent interactions which are responsible for the fine structure splitting of the P-wave triplet and can also shift the center-of-gravity mass.
V Analysis of the Hypothetical Toponium System
The successful application of the bootstrap method to charmonium and bottomonium motivates its extension to the heaviest quarkonium system: toponium ($t\bar{t}$). This serves as a valuable theoretical exercise to probe the behavior of the non-relativistic quarkonium potential and the bootstrap method at higher mass scales.
A crucial distinction must be made at the outset. Unlike charmonium and bottomonium, a stable, long-lived toponium meson is not expected to exist in nature. This is due to a fundamental race between the timescales of the weak and strong interactions. The top quark, with its enormous mass of $m_{t}\approx 172.52$ GeV [10], decays via the electroweak interaction ($t\to Wb$) with an exceptionally short mean lifetime of $\tau_{\text{decay}}\approx 5\times 10^{-25}$ s [10]. In contrast, the characteristic timescale for the strong interaction to bind quarks into a hadron, the hadronization time, is approximately $\tau_{\text{had}}\approx 1/\Lambda_{\text{QCD}}\sim 3\times 10^{-24}$ s [2, 10].
Since $\tau_{\text{decay}}\ll\tau_{\text{had}}$, the top quark decays, on average, an order of magnitude faster than the strong force can form a conventional bound state [11, 12]. This makes the top quark the only "bare" quark that can be studied before it is obscured by the non-perturbative dynamics of hadronization.
Despite its physical impossibility, a theoretical calculation for a hypothetical stable toponium ground state ($1S$, $l=0$) can be performed using the quantum bootstrap method. This provides a clean test of the self-consistency of the non-relativistic quarkonium potential when extrapolated to the high-mass regime. Using the potential parameters $\alpha_{s}=0.1088$, $B=0.18$ and a top quark mass of $m_{t}=172.5$ GeV, the bootstrap calculation yields a stable and consistent prediction for the ground state properties, summarized in Table 2.
Table 2: Predicted ground state properties of a hypothetical stable toponium ($1S$) state using the quantum bootstrap method.
Quantity Predicted Value
Binding Energy ($E_{1S}$) -0.6855 GeV
Total Mass ($M=2m_{t}+E_{1S}$) 344.3145 GeV
The most striking feature of this result is the large negative binding energy of approximately -0.69 GeV, which is substantially larger in magnitude than those of charmonium ($\sim 0.8$ GeV) and bottomonium ($\sim 0.5$ GeV). This result is a direct and physically intuitive consequence of structure of the non-relativistic quarkonium potential. The much heavier top quarks would form an extremely compact system with a very small average inter-quark distance $\langle r\rangle$. At these very short distances, the attractive Coulomb-like term (A/r-A/r) of the potential becomes overwhelmingly dominant over the linear confinement term ($Br$). This strong, short-range attraction leads to a very tightly bound state, which is reflected in the large negative binding energy.
The theoretical exercise of calculating the properties of a hypothetical toponium state can be of significance when juxtaposed with recent experimental breakthroughs at the Large Hadron Collider. While a stable toponium meson is precluded, non-relativistic QCD (NRQCD) predicts that the strong interaction should still manifest as a "quasi-bound state"—a broad, short-lived resonance-like structure just below the $t\bar{t}$ production threshold. The experimental signature of such a state would not be a sharp mass peak, but a distinct enhancement in the production cross-section near $m_{t\bar{t}}\approx 2m_{t}$.
In a landmark development in 2025, both the CMS and ATLAS collaborations reported the observation of just such a phenomenon [11, 12].
The energy at which this enhancement is observed, near the production threshold of $m_{t\bar{t}}\approx 345$ GeV, is remarkably close to the bootstrap prediction of $M\approx 344.3$ GeV for the hypothetical ground state mass (Table 2). This agreement between the experimental result and the prediction from a well-motivated phenomenological potential is a powerful validation of the entire physical picture. Furthermore, it showcases the quantum bootstrap as a reliable tool for extracting physical observables from such a potential.
This calculation provides a compelling theoretical counterpart to the experimental observation of the toponium quasi-bound state. While the experiments probe the dynamic formation of a short-lived resonance, our bootstrap calculation solves for the static properties of its would-be ground state. The proximity of the results suggests that the energy of this transient state is indeed governed by the same underlying potential that describes stable quarkonia.
VI Conclusion
In this paper, the quantum bootstrap method was successfully applied to the problem of heavy quarkonium spectroscopy as described by a non-relativistic potential. By leveraging the fundamental principles of quantum mechanics, our approach provides a robust, non-perturbative, and computationally efficient alternative to traditional methods for solving the Schrödinger equation. This work represents a successful synergy between a non-perturbative computational method and experimental particle physics.
A primary extension of this work would be the inclusion of spin-dependent interactions. The current model calculates the spin-averaged "gross structure" of the quarkonium spectrum. By incorporating relativistic corrections via the Breit-Fermi Hamiltonian [7], the fine and hyperfine structure can be modelled. This would involve adding spin-orbit, tensor, and spin-spin interaction terms to the potential. A bootstrap analysis of this more complex Hamiltonian could predict the mass splittings within the P-wave triplets ($\chi_{c},\chi_{b}$) and, crucially, distinguish between the pseudoscalar ($1^{1}S_{0}$, e.g., $\eta_{c}$) and vector ($1^{3}S_{1}$, e.g., $J/\psi$) ground states.
The calculations for the spin-averaged mass spectra of the charmonium and bottomonium systems are in good agreement with experimental data, with accuracies exceeding 99.9% for the ground states. This result validates this bootstrap approach as a high-precision tool for quantum mechanical problems.
Finally, this bootstrap calculation is used to predict the ground state mass of a hypothetical toponium state, obtaining a value of $M\approx 344.3$ GeV, in agreement with the observed enhancement at $m_{t\bar{t}}\approx 345$ GeV by the ATLAS and CMS experiments. This result highlights the potential of the quantum bootstrap as a method capable of providing meaningful theoretical guidance to frontier experiments. This work bridges a purely theoretical calculation with a landmark discovery, offering a new perspective on the dynamics of the heaviest quark and reinforcing the deep connections between fundamental principles and observable reality.

### v3

@@ 4.1 Bootstrap spectra for heavy quarkonia
The results show the effectiveness of the quantum bootstrap method. For the ground states ($l=0$), the predicted masses for $J/\psi$ and $\Upsilon$ are accurate, with deviations of less than 2 MeV and 6 MeV.The predictions for the P-wave states ($l=1$) are also close, with deviations of less than 7 MeV. This slight deviation is expected, as this model calculates the spin-averaged mass and does not include spin-dependent interactions which are responsible for the fine structure splitting of the P-wave triplet and can also shift the center-of-gravity mass.
Traditional potential models ([13, 14, 16]) solve the Schrödinger or spin-dependent Dirac equations with fitted parameters $(\alpha_{s},\kappa,m_{Q})$, while Lattice NRQCD simulations [17, 18] compute spectra from first principles at the cost of large-scale computation. The typical accuracies achieved by these methods are comparable with the one obtained with the bootstrap method (below $1\%$). However, this method relies solely on moment consistency and positivity, without solving a differential equation.
In contrast with the Coulomb potential, which introduces a singularity at $r=0$, making the corresponding Stieltjes moment problem ill-conditioned for low truncation depth, the potential in equation 1 introduces a linear term that regularizes the recurrence structure and improves the conditioning of the Hankel matrices.
Moreover, the physical wavefunctions for $\ell\geq 1$ vanish at $r=0$, avoiding the divergence that limits the bootstrap precision in purely Coulombic systems. These features explain the rapid and stable convergence of the bootstrap for heavy quarkonia.
Applying the same framework to the top–antitop system yields a spin-averaged $1S$ centroid at approximately $344~\mathrm{GeV}$. This value lies close to the $t\bar{t}$ production threshold and is consistent with the threshold enhancement recently observed by the CMS and ATLAS collaborations [11, 12]. Although the top quark decays before hadronization, non-relativistic QCD predicts Coulombic enhancements below and near threshold [19, 20], and recent QCD calculations including bound-state effects corroborate such behaviour [21]. Accordingly, the bootstrap result should be regarded as a theoretically motivated centroid consistent with the observed threshold behaviour.
These results place the quantum bootstrap method as a promising approach to heavy-quark bound states, on equal footing with other non-perturbative approaches. While lattice QCD provides first-principles validation, the bootstrap method captures the essential spectral structure from purely algebraic constraints. This complementarity suggests that bootstrap-inspired positivity frameworks could eventually serve as analytic surrogates for certain lattice calculations, closing the gap between non-relativistic quantum mechanics and quantum field theory.
5 Conclusion
We have presented a non-perturbative determination of quarkonium spectra using the quantum-mechanical bootstrap approach. Instead of solving the Schrödinger equation explicitly, the method reconstructs the allowed energies from algebraic moment recursions and the positivity of Hankel matrices. Applied to a non-relativistic quarkonium potential 1, the bootstrap method reproduces the spin-averaged $1S$ and $1P$ levels of charmonium and bottomonium and predicts a quasi-bound $1S$ centroid near $344~\text{GeV}$ for a hypothetical toponium system. These results demonstrate that spectral information can be extracted purely from algebraic consistency and positivity, without specifying any functional basis or boundary conditions.
Beyond its numerical accuracy, the bootstrap provides a conceptual link between formal positivity conditions in quantum field theory and concrete bound-state spectroscopy in quantum mechanics. The success of the method highlights that the rate of convergence is governed by the analytic structure of the potential and by the conditioning of the associated moment problem. Compared with conventional potential models or lattice QCD, the bootstrap is computationally inexpensive, scalable, and naturally suited for systematic convergence analysis.
Future work can extend this framework in several directions. Including spin-dependent and relativistic corrections would allow a full description of fine and hyperfine splittings, while applying the bootstrap to higher-dimensional or field-theoretical systems—such as two-particle correlation functions or effective Hamiltonians in QCD—could provide a new analytical window into non-perturbative dynamics. Overall, the results reported here support the view that positivity-based bootstrap methods offer a plausible route to the computation of bound-state spectra across quantum mechanics and quantum field theory.